can you help me
solve by factoring and using the principle of zero products
x^2+4x-77=0

Answer by anantha(86)

sol:
x^2+4x-77=0
product of two numbers we get 77
and adding or subtracting we get 4
those numbers are 11 and 7
their product=11*7=77
subtracting those numbers=11-7=4
above equation can be written as
x^2+11x-7x-77=0
here 4 terms are there ,in 1 and 2 terms x is common
and 3 and 4 terms 7 is common
x(x+11)-7(x+11)=0
now x+11 is common term
(x+11)(x-7)=0
x+11=0
x=-11
x-7=0
x=7
the roots are 7 and -11